OpenStudy (anonymous):

∠DEF and ∠ABC are supplementary angles and ∠DEF is four times as large as ∠ABC. Determine the measure of each angle. pls show wrk

OpenStudy (anonymous):

supp. means they add up to 180. 180/5 =36 def= 4*36 def= 144 abc= 36

OpenStudy (anonymous):

thanx that makes sense